### v2

# TimeFlow: Longitudinal Brain Image Registration and Aging Progression Analysis

Abstract
Predicting future brain states is crucial for understanding healthy aging and neurodegenerative diseases. Longitudinal brain MRI registration, a cornerstone for such analyses, has long been limited by its inability to forecast future developments, reliance on extensive dense longitudinal data, and the need to balance registration accuracy with temporal smoothness. In this work, we present TimeFlow, a novel framework for longitudinal brain MRI registration that overcomes all these challenges. TimeFlow leverages a U-Net architecture with temporal conditioning inspired by diffusion models, enabling accurate registration using only two images as input and facilitating prospective analyses through future image prediction. Unlike traditional methods, TimeFlow eliminates the demand for explicit smoothness regularizers and dense sequential data while maintaining temporal consistency and continuity. Experimental results highlight its superior performance in both future timepoint prediction and registration accuracy compared to state-of-the-art methods. Additionally, TimeFlow supports novel biological brain aging analyses, effectively differentiating neurodegenerative conditions from healthy aging, all without requiring segmentation thus avoiding non-trivial annotation and inconsistent segmentation flaws. This framework paves the way for accurate, data-efficient, and annotation-free prospective analyses of brain aging and chronic diseases.
Figure 1: Outstanding performance of TimeFlow for future brain MRI prediction. TimeFlow demonstrates the superior capability of future extrapolation compared to (i) NODER, which employs NeuralODE for geodesic regression, and (ii) VoxelMorph with linear scaling. Prediction errors in these methods are marked with red arrows. Given two previous timepoints (Baseline and Year 3), TimeFlow can extrapolate precise and smooth deformation to predict brain MRI in Year 6.
1 Introduction
Understanding how the human brain changes over time — both in the course of healthy aging and under the influence of neurodegenerative diseases — remains a central challenge in neurology and related fields [33, 34]. Such insights not only shed light on the underlying pathological processes but also pave the way for more individualized treatment planning and earlier, more targeted interventions [6]. Magnetic Resonance (MR) imaging has become the standard modality for capturing these longitudinal changes, owing to its high spatial resolution and excellent soft-tissue contrast. By repeatedly scanning the same individual, clinicians and researchers can visualize evolving patterns of atrophy, lesion progression, and other structural transformations as they unfold [21, 2].
Despite advances in longitudinal registration [17, 43], which aligns a subject’s brain images across multiple time points, two critical challenges persist: 1. Predicting future brain states and analyzing aging progression is one of the ultimate goals of longitudinal registration. While retrospective assessments of past changes are valuable, the ability to predict future brain transformations would dramatically enhance clinical utility [27] and enable clinicians to forecast atrophy trajectories and guide more effective interventions [23]. By extrapolating physically plausible transformations, longitudinal registration enables future scan prediction through simple image warping, avoiding the need for complex generative approaches. Moreover, comparing the predicted future brain changes with the actual observed brain state allows clinicians to assess whether the disease is advancing at an accelerated rate compared to historical progression. This insight can provide early warnings of disease exacerbation and help refine patient-specific prognostic models. Recent works [50, 3, 51] primarily focus on modeling the observed sequence of scans, allowing interpolation between observed timepoints but lacking the capability to extrapolate beyond the final scan, as illustrated in Fig. 1. This limitation makes it difficult to assess deviations from expected progression, anticipate disease worsening, and adjust treatment accordingly. 2. Ensuring temporal continuity with limited sequential data. Existing methods impose temporal smoothness constraints [49, 53, 9] and often require densely sampled longitudinal data (often more than five scans) [28, 36], which is impractical in typical clinical scenarios where only a few follow-up scans (<3) are available.
We introduce TimeFlow, a novel longitudinal brain MRI registration capable of predicting future brain changes. The key contribution is a novel constraint that ensures temporal consistency and continuity — removing the demand for smoothness regularizers. Moreover, TimeFlow can operate longitudinal registration with only paired images in inference, eliminating the requirement for long image sequences. Finally, TimeFlow can extrapolate to future time points and analyze aging progression without relying on segmentation masks. Our contributions are twofold:
  1. 1.
    Inspired by conditioning mechanisms used in diffusion models, TimeFlow employs a U-Net architecture conditioned on a temporal parameter reflecting biological aging. This temporal conditioning allows for generating temporally continuous and smooth deformation fields, even when only two images are used as input. Moreover, this approach enables future brain images prediction, surpassing state-of-the-art methods in accuracy and smoothness.
  2. 2.
    We develop a novel methodology for longitudinal brain aging and disease progression analysis, utilizing TimeFlow’s advanced predictive capabilities. Our approach quantifies relative aging progression using healthy subjects as a reference baseline, where biological aging is assumed to align with chronological aging. This method enables the identification and differentiation of subjects with Mild Cognitive Impairment (MCI) and Dementia, who typically exhibit accelerated biological aging compared to chronological aging. Remarkably, our method performs this analysis solely through registration, eliminating the need for brain structure annotations and avoiding potential segmentation-related error propagation.
-sho
2 Related Work
Cross-sectional registration
methods laid the groundwork for medical image registration, pioneering this field with techniques such as parametric spline-based models [41, 29], optical flow [35, 52], and diffeomorphic transformations [44, 47, 5]. However, they demand prohibitive computing time in the order of several hours to register one single subject, posing a significant barrier to clinical deployment. To address this, learning-based frameworks [4, 31, 37, 7, 14, 22] have emerged, dramatically reducing inference time and often improving registration accuracy. Nevertheless, all such methods rely on carefully tuned smoothness regularizers to ensure anatomically plausible deformation fields. An improper weighting factor of smoothness regularizers can either reduce fidelity or introduce unrealistic warps. Additional tuning or adaptive strategies are often needed to strike an optimal balance [18, 32, 48].
Longitudinal registration
methods [42, 19, 20, 11, 12, 36, 9, 24] often leverage the backbone architecture of the cross-sectional registration, whereas one additional constraint is introduced: the temporal continuity/consistency across multiple timepoints of a subject. Two major strategies are often used to tackle this problem. First, an explicit smoothness constraint along the temporal dimension is incorporated to maintain coherent deformation fields over time [53, 28]. However, as with spatial smoothness, temporal regularization faces a similar trade-off: enforcing temporal continuity can inadvertently lower spatial accuracy. Second, many longitudinal methods require multiple, densely sampled scans (>10) to capture temporal dynamics [3, 50]. In real-world clinical scenarios, patients often have only a few follow-up scans ($\leq 3$), restricting model flexibility and practical implementation. Importantly, these methods commonly can only carry out retrospective analyses and are not capable of prospective studies such as future image extrapolation and aging/disease progression risk prediction.
3 Methods
Figure 2: Left: Architecture of TimeFlow. TimeFlow $f_{\theta}$ integrates time $t$ information via adaptive instance normalization. The network takes the time $t$, the images at first $I_{0}$ and last visit $I_{L}$ as input, and predicts the temporal deformations. Examples of $t=0.5$ (interpolation at middle), $t=1$ (endpoint registration), and $t=2$ (extrapolation) are shown. Right: Interpolation (upper yellow) and Extrapolation (bottom green) Consistency. TimeFlow can predict bidirectional time-continuous deformation fields. The symmetric midpoint similarity forces the deformed images to be aligned at time $t_{k}$, while the midpoint interpolation flow loss encourages the deformation composition consistency. The extrapolation similarity and flow enforcement give information on what a realistic and reasonable extrapolation should be. Training with both consistency losses enables TimeFlow to generate plausible deformation in $t\in[-\infty,+\infty]$. The input images are highlighted by blue squares (Input), while the supervision are highlighted by dot-lined squares (Sup.).
3.1 Time-Continuous Registration
Given a sequence of longitudinal images $\{I_{k}\mid k=0,\ldots,L\}$, we aim to find a temporally continuous deformation trajectory that connects all the discrete time points. To achieve this, we construct a time-continuous registration function $\phi_{t}=f_{\theta}(I_{0},I_{L},t)$. It takes the first visit $I_{0}$ and the last visit $I_{L}$ as input images and predicts the deformation $\phi_{t}$ at time $t$. As shown in Figure 2, we utilize a time-embedded UNet, frequently used in diffusion models [15, 40] to learn the mapping function. The time variable $t$ is first projected to the latent embedding space by a sinusoidal positional encoding layer and two-layer MLPs. Next, the time latent vector is fed to two separate MLPs to steer the mean and variance of the inter-layer adaptive instance normalization. The network is thus instilled with the capability of generating meaningful time embedding and predicting time-continuous flow fields by the novel time conditioning mechanism. Moreover, to ensure the predicted deformation trajectory passes through each visit $I_{k}$ at time $t=t_{k}$, we incorporate the intermediate time points as supervision in training.
3.2 Interpolation Constraints
Intermediate Similarity.
Our goal is to learn a time-continuous registration function $\phi_{t}^{0\rightarrow L}=f_{\theta}(I_{0},I_{L},t)$. When $t=0$, we have $f_{\theta}(I_{0},I_{L},0)=\textrm{Id}$, the identity mapping. When $t\in(0,1]$, $\phi_{t}^{0\rightarrow L}$ is the forward deformation mapping from $I_{0}$ towards $I_{L}$, and $\phi_{1}^{0\rightarrow L}$ moves $I_{0}$ to reach $I_{L}$. Similarly, for negative values, $t<0$, $\phi_{t}$ maps backward from $I_{L}$ to $I_{0}$. Exploiting bidirectional time-continuous deformations, we can formulate a symmetric similarity measure:
$\mathcal{L}_{\textrm{sim}}(I_{0}\circ\phi_{t_{k}}^{0\rightarrow L},I_{k})+\mathcal{L}_{\textrm{sim}}(I_{L}\circ\phi_{t_{k}-1}^{0\rightarrow L},I_{k}),$ (1)
which deforms both images to remove the bias introduced by single-side warping [1, 30, 16]. As shown in the yellow part in Fig. 2(b), $I_{0}\circ\phi_{t_{k}}^{0\rightarrow L}$ and $I_{L}\circ\phi_{t_{k}-1}^{0\rightarrow L}$ describe the deformation of both $I_{0}$ and $I_{L}$ to the time $t=t_{k}$. Given an acquired MRI $I_{k}$ at time $t=t_{k}$, we can simply use it to compute the symmetric intermediate similarity and thus force the network to learn plausible deformation trajectories from $I_{0}$ to $I_{L}$ that pass through intermediate images.
Generalized Intermediate Similarity.
However, in real-world clinical scenarios, many subjects will only have MRI from two time points available. In such a case, we rewrite the symmetric similarity as:
$\mathcal{L}_{\textrm{sim}}(I_{0}\circ\phi_{\hat{t}}^{0\rightarrow L},I_{L}\circ\phi_{\hat{t}-1}^{0\rightarrow L}),$ (2)
where $\hat{t}$ is sampled from a uniform distribution $\mathcal{U}(0,1)$. The registration is guided properly through this synthetic variant of the symmetric similarity, even without any intermediate observation as supervision. Moreover, taking advantage of the freedom to choose $\hat{t}$, the network can learn densely on the time axis and is no longer constrained to intermediate visits at specific times in the dataset.
Extrapolation Interpolation
[…]
Seq-VXM - 4.5$\pm$0.9 25.2$\pm$1.7 2.6$\pm$0.3 0.0$\pm$0.1
(Upper bound)
rigid 10.3$\pm$1.9 16.9$\pm$1.5 - - 7.3$\pm$1.1 20.6$\pm$1.5 - -
NODER 9.3$\pm$1.9 18.0$\pm$1.7 1.8$\pm$0.9 0.2$\pm$1.9 5.5$\pm$0.9 23.2$\pm$1.5 1.8$\pm$0.5 0.0$\pm$0.3
VXM 9.1$\pm$1.9 18.3$\pm$1.9 6.3$\pm$4.8 $10^{4}\pm 10^{4}$ 5.6$\pm$1.1 22.9$\pm$1.7 1.5$\pm$0.2 0.2$\pm$0.4
VXM-diff 9.2$\pm$2.0 18.4$\pm$1.9 6.3$\pm$2.0 $10^{3}\pm 10^{3}$ 5.5$\pm$1.0 23.2$\pm$1.7 2.2$\pm$0.3 0.1$\pm$0.2
TimeFlow 8.6$\pm$1.6 18.8$\pm$1.7 5.0$\pm$1.0 $10^{3}\pm 10^{3}$ 5.5$\pm$1.0 23.0$\pm$1.7 1.6$\pm$0.1 0.1$\pm$0.4
TimeFlow-diff 8.8$\pm$1.8 18.7$\pm$1.9 6.0$\pm$1.0 6.1$\pm$11.2 5.3$\pm$1.0 23.3$\pm$1.8 2.2$\pm$0.2 0.0$\pm$0.0
[…]
Interpolation Consistency To encourage plausible time-continuous deformations, we propose the interpolation consistency as a regularizer:
$\begin{split}\mathcal{L}_{\textrm{inter-consis-flow}}=&\|\phi_{1}^{0\rightarrow L}\circ\phi_{t_{k}-1}^{0\rightarrow L}-\phi_{t_{k}}^{0\rightarrow L}\|_{2}^{2}\\ +&\|\phi_{-1}^{0\rightarrow L}\circ\phi_{t_{k}}^{0\rightarrow L}-\phi_{t_{k}-1}^{0\rightarrow L}\|_{2}^{2},\end{split}$ (3)
which combines the two constraints shown in Fig. 2(b). We formulated the loss with the notation $t_{k}$, but as described in the previous paragraph, $t_{k}$ can also be sampled from the uniform distribution if no intermediate observation exists. It is derived from the common transformation composition consistency that two consecutive transforms are equivalent to a direct transform. Moving from $t=0$ to $t=1$ and then back to $t=t_{k}$ is equivalent to moving directly from $t=0$ to $t=t_{k}$ (arrow 1 composed with arrow 3 should equal to arrow 2 in Fig. 2(b)). The backward case is derived similarily. It is worth noting that this differs from the commonly used smoothness regularizations, which explicitly enforce a smooth gradient over the flow field. Therefore, there is only a minimal trade-off between accuracy and smoothness since both the similarity and the flow consistency encourage plausible and reasonable longitudinal transformations.
3.3 Extrapolation Consistency
Through the intermediate similarity and interpolation consistency, we obtain meaningful predictions for bidirectional registrations in $t\in[-1,1]$. As our main objective is extrapolation, we formulate criteria for the intervals $[1,+\infty)$ (forward) and $(-\infty,-1]$ (backward) next. To this end, we consider the deformation $\phi_{1}^{0\rightarrow k}$ from $I_{0}$ to $I_{k}$. Given the time of $I_{k}$ being $t_{k}$, it holds that $\phi_{t_{k}}^{0\rightarrow L}=\phi_{1}^{0\rightarrow k}$. As illustratedin Fig. 2(b), they both register $I_{0}$ to $I_{k}$. Consequently, we can then formulate $\phi_{1}^{0\rightarrow L}=\phi_{1/t_{k}}^{0\rightarrow k}$ (arrow 1 equals to arrow 5 in Fig. 2(b)), where $\phi_{1/t_{k}}^{0\rightarrow k}$ denotes an extrapolation because $t_{k}\in(0,1)$ and therefore $11/t_{k}>$ 1. Based on this notation, we define the extrapolation similarity as:
$\mathcal{L}_{\textrm{sim}}(I_{0}\circ\phi_{1/t_{k}}^{0\rightarrow k},I_{L}).$ (4)
Again, we can either use intermediate observations with time $t_{k}$ or sample from $\mathcal{U}(0,1)$. If $t_{k}$ is sampled from $\mathcal{U}(0,1)$, i.e., $I_{k}$ is not observed but only $I_{0}$ and $I_{L}$, we then define $I_{k}=I_{0}\circ\phi_{t_{k}}^{0\rightarrow L}$. In addition to the extrapolation similarity measure, we define the forward extrapolation flow consistency as:
$\mathcal{L}_{\textrm{ext-consis-flow}}=\|\phi_{1}^{0\rightarrow L}-\phi_{1/t_{k}}^{0\rightarrow k}\|_{2}^{2}.$ (5)
As shown in the bottom green part of Fig. 2(b), both $\phi_{1}^{0\rightarrow L}$ and $\phi_{1/t_{k}}^{0\rightarrow k}$ map from $I_{0}$ to $I_{L}$, thus should be consistent with each other. The backward versions of the extrapolation similarity and extrapolation flow consistency are:
$\begin{split}&\mathcal{L}_{\textrm{sim}}(I_{L}\circ\phi_{1/(t_{k}-1)}^{k\rightarrow L},I_{0})\quad\text{and}\\ &\mathcal{L}_{\textrm{ext-consis-flow}}=\|\phi_{-1}^{0\rightarrow L}-\phi_{1/(t_{k}-1)}^{k\rightarrow L}\|_{2}^{2}.\end{split}$ (6)
The similarity measures and flow consistencies both contribute to achieving a realistic and coherent extrapolation. The proposed losses effectively ensure accurate extrapolation while implicitly capturing temporal consistency and continuity, without the need for additional temporal smoothness constraints.
4 Data and Experiments
Figure 3: Visualization of Interpolation and Extrapolation Results. TimeFlow/diff takes the Baseline and Year 4 images as input, interpolates to Year 1.5, and extrapolates to Year 8. The first row displays the ground truth images. The second row shows the moved baseline image. The third and fourth rows display error maps: the initial error map before registration and the error map after TimeFlow/diff registration. Error values range from $[-1,1]$ due to the image intensity normalization to $[0,1]$. These results demonstrates that TimeFlow/diff achieves high accuracy in both temporal interpolation and extrapolation. Detailed temporal deformations can be referred to Appendix.
Data and Pre-processing.
We evaluate our method on the public Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset, which comprises longitudinal T1-weighted Magnetic Resonance (MR) scans from both healthy subjects and patients diagnosed with Alzheimer’s Disease (AD) or MCI, i.e., neurodegenerative disorders that exhibit specific patterns of atrophy development in the brain. We employ the longitudinal processing from FreeSurfer v7.2 [13]. First, all scans are processed individually, including skull stripping, normalization. Subsequently, the image sequence is rigidly aligned via a groupwise rigid registration [39]. FreeSurfer aggregates information across time points and refines the pre-processed outputs. We first select 134 subjects who underwent substantial anatomical change (median change magnitude within the brain is $>0.3$mm) over time. Then, the dataset is divided into training, validation, and testing sets, ensuring a balance in diagnosis, age, sex, and the number of visits. It results in 74, 30, 30 subjects for each split, respectively, with the number of time points per subject ranging from 2 to 12. All input scans have isotropic 1mm spacing and are cropped to a uniform size of $160\times 160\times 192$ voxels. The intensity values are scaled to the range [0,1], from percentile [0,99.9] without clipping.
Implementation Details.
Our model is implemented using Pytorch 2.1.2. and training/test is carried out on one NVIDIA A100 40G with a batch size of $1$. We set the dimension of time embedding to $16$, the feature channels in the U-Net encoder and decoder to $[32,32,48,48,96]$. In addition to TimeFlow, which outputs the time-dependent deformation field directly, we have the diffeomorphic variant TimeFlowdiff, which utilizes the scaling and squaring layer with $7$ integration steps to integrate the temporal stationary velocity field. The weights of both interpolation and extrapolation similarity are set to $1.0$. The weights of interpolation and extrapolation flow consistency are set to $2,0.03$ for TimeFlow and $1.25,0.025$ for TimeFlowdiff.
Baselines.
To evaluate the registration performance we compare TimeFlow with the following methods:
  • NODER [51, 3], which is aware of every time point and takes the complete image sequence as the model input. It carries out subject-specific optimization and models the time-varying velocity field with a neural network and leverages Neural ordinary differential equation (ODE) [8] to compute integration and adjoint gradients. It fits the subject-specific temporal deformation trajectory by maximizing the similarity with each frame.
  • VoxelMorph (VXM) [4] and the extended diffeomorphic VoxelMorph (VXMdiff[10]. Since both are pairwise registration methods, the interpolation and extrapolation prediction is carried out based on the linear scaling of the displacement or velocity field from the pairwise registration. The integration steps of VXMdiff is also set to 7.
  • Sequence-based VoxelMorph (Seq-VXM) [50]. Seq-VXM processes the whole sequence and registers the baseline image to each intermediate time point. Given that Seq-VXM performs multiple pairwise registrations by taking the whole image sequence as model input, it is aware of all images at every time point. Therefore, we treat Seq-VXM as the performance upper bound for the retrospective study (refer to Tab. 1).
Figure 4: Visualization of the Extrapolation Results. Left: Error maps show differences between the extrapolated predictions and the ground truth future images. Right: Future image predictions for different $t$ values, where no ground truth is available. $t=1$ represents one-step pairwise registration between $I_{m}$ and $I_{f}$. Thus, $t=2,4,6$ represents two, four, six-step extrapolation over $I_{f}$. Major errors are highlighted with red arrows. VXM/diff produces unsmooth deformations, leading to structural inconsistencies and loss of anatomical details (zoom in for more details).
For all learning-based methods, we apply Local Normalized Cross Correlation (LNCC) as the training loss, the same as in TimeFlow/diff. The radius of LNCC in NODER is set to $r=9$ while for the other methods, it is set to $r=4$. The loss function weighting factors in NODER follows their original suggestion. Seq-VXM, VXM and VXMdiff compute the SOTA smoothness regularizer GradICON [45, 46] with weight $1.5,0.5,1.0$.
Evaluation metrics.
[…]
Figure 5: Left: Progression Analysis for Control (CN), MCI and Dementia. A progression rate closer to 1 indicates slower pathological brain structural change. Right: Relative Progression of 9 representative subjects. Each colored dot represents one visit. The $x$-axis denotes the years since the first visit. The $y$-axis stands for the relative progression rate compared to the baseline visit. A higher value indicates faster brain structural changes. The grey line in the background is the mean fitted linear line across subjects. Colormap: CN-green, MCI-orange, Dementia-red.
5 Results
5.1 Prospective studies: Predicting the future
We first evaluate TimeFlow’s performance in prospective studies, i.e., for extrapolating the deformations to predict follow-up images. Our model can perform this prospective analysis in two ways: 1. Based on two available scans and given a future timepoint $t$ ($t>1$), our model can directly predict the extrapolated flow field and the warped image at $t$. 2. Given a future scan, our model can estimate the corresponding $\phi_{t}$, thus the time $t$, which can reflect the subjects’ biological aging progression.
Given t, predict the future image.
As discussed in Sec. 3, one of TimeFlow’s biggest benefits is its ability to naturally extrapolate the deformation and predict a future image given $t>1$. The quantitative evaluation of extrapolation is showcased in table Tab. 1. TimeFlow demonstrates superior extrapolation accuracy compared to all baselines with the lowest MAE and the highest PSNR. Seq-VXM is not compared here since it cannot carry out future predictions. Fig. 3 (Year 8) and Fig. 4 (Left) illustrate TimeFlow’s extrapolation prediction when ground truth images are available, while Fig. 4 (Right) demonstrates the extrapolation given a large $t$ where ground truth is not available anymore. As shown in Fig. 4, NODER extrapolates unrealistic deformations, leading to significant error in future image predictions. VXM/diff, which applies linear scaling to flow fields, results in unsmooth deformation and dissolves the brain, causing anatomical details to fade. In contrast, TimeFlow/diff generates smooth and accurate extrapolated flow fields, aligning with the lowest MAE and low NDV in Tab. 1.
Given future scan, assess the brain aging.
We can utilize the extrapolation capability of TimeFlow for disease progression analysis. For each future scan, we predict a corresponding $t$ value based on prior visits. We compare the relative brain aging rate across diagnosis groups in the ADNI dataset, i.e., Controls, MCI, and Dementia. Specifically, MRI triplets are sampled from the validation and test set, where the intervals between the first and second timepoints are similar to the intervals between the second and third timepoints. The first two time points are used as input to extrapolate to the third, future time point. By comparing with the ground truth MRI, we compute the optimal extrapolated time $t_{\textrm{ext}}$. From this, we can compute the brain aging rate by dividing the extrapolated time by the actual time interval. The idea is that if brain aging is accelerated, i.e., in MCI or Dementia patients, the progression rate will be larger than 1, meaning that brain aging is faster than chronological aging. As shown in the boxplot of Fig. 5, Dementia exhibits the highest progression rate, followed by MCI, while the control group remains is close to $1$, highlighting a biologically plausible separation of groups based with our brain aging rate metric. To visualize the brain aging rate over time, we plot subject-specific progression curves for 9 representative subjects (3 from each group). We observe accelerated progression compareed to the baseline visit in MCI and Dementia, while the control group (CN) maintains a flatter curve. This highlights how TimeFlow’s extrapolation capabilities enables meaningful prospective analysis on Alzheimer’s disease, without the need for additional anatomical annotations.
$\mathcal{L}_{\textrm{inter}}$ $\mathcal{L}_{\textrm{ext}}$ Extrapolation Interpolation
[…]
1.58 $\pm$ 0.17 16.7$\pm$7.2 13.9$\pm$4.1 $>10^{3}$ $>10^{7}$ 5.3$\pm$1.0 23.3$\pm$1.7 4.7$\pm$1.3 $>10^{4}$
0.82 $\pm$ 0.18 9.8$\pm$2.2 17.6$\pm$2.1 9.7$\pm$4.8 $>10^{4}$ 5.5$\pm$1.0 23.0$\pm$1.7 1.8$\pm$0.1 0.3$\pm$0.7
0.62 $\pm$ 0.12 8.7$\pm$1.7 18.7$\pm$1.8 5.3$\pm$1.4 $>10^{3}$ 5.5$\pm$1.0 23.0$\pm$1.7 1.6$\pm$0.1 0.1$\pm$0.4
0.60 $\pm$ 0.11 8.6$\pm$1.6 18.8$\pm$1.7 5.0$\pm$1.0 $10^{3}\pm 10^{3}$ 5.5$\pm$1.0 23.0$\pm$1.7 1.6$\pm$0.1 0.1$\pm$0.4
3.25 $\pm$ 0.25 27.5$\pm$16.2 11.4$\pm$6.2 $>10^{2}$ $>10^{6}$ 5.5$\pm$1.1 22.8$\pm$1.9 6.9$\pm$2.2 $>10^{2}$
0.82 $\pm$ 0.12 10.4$\pm$2.9 17.2$\pm$2.7 8.4$\pm$2.7 $>10^{3}$ 5.3$\pm$1.0 23.4$\pm$1.8 2.5$\pm$0.2 0.3$\pm$0.7
0.55 $\pm$ 0.09 8.8$\pm$1.8 18.6$\pm$1.9 6.1$\pm$1.1 18.8$\pm$39.4 5.3$\pm$1.0 23.3$\pm$1.8 2.3$\pm$0.2 0.0$\pm$0.1
0.56 $\pm$ 0.09 8.8$\pm$1.8 18.7$\pm$1.9 6.0$\pm$1.0 6.1$\pm$11.2 5.3$\pm$1.0 23.3$\pm$1.8 2.2$\pm$0.2 0.0$\pm$0.0
[…]
5.2 Retrospective studies: Matching the deformation trajectory
We conduct studies on timepoints interpolation to assess whether TimeFlow learns temporally continuous and plausible deformation field connecting intermediate visits. In this setting, the baseline visit $I_{0}$ and the last scan $I_{L}$ serve as inputs, and the deformation fields $\phi_{t_{k}}$ are estimated to register intermediate images $I_{k}$, where $0<K<L$. Quantitative comparisons with baseline methods are presented in Tab. 1. As discussed in Sec. 4, Seq-VXM process the entire sequence and performs explicit pairwise registration for each timepoint, making it the performance upper bound. Despite using only the baseline and last scan as inputs, TimeFlow outperforms NODER, which is also aware of the entire sequence and parametrized by NeuralODE. Additionally, TimeFlowdiff achieves the least non-diffeomorphic volume without any extra spatial smoothness regularization. The interpolation results shown in Fig. 3 (Year 1.5) futher demonstrate that TimeFlow achieves almost zero residual errors compared to the initial misalignment.
5.3 Ablation Study
As described in Sec. 3 and supported by both quantitative and qualitative results, TimeFlow achieves temporally continuous and spatially smooth deformation fields without requiring explicit temporal or spatial smoothness regularization. To assess the effectiveness of our proposed consistency losses, we conduct ablation studies. To evaluate the temporal smoothness and continuity of TimeFlow, we compute the derivative $\partial\phi_{t}/\partial t$ using the finite difference method by sampling $t$ at intervals of $\Delta t=0.25$. Similar to the commonly used diffusion regularizer in learning-based registration methods, which computes the Dirichlet energy over spatial gradients, we quantify temporal smoothness as $\|\partial\phi_{t}/\partial t\|$. Specifically, we compute the norm of the temporal displacement derivatives within the brain and then take the mean over all brain voxels as the final metric. As shown in Tab. 2, incorporating interpolation consistency loss significantly improves both temporal and spatial smoothness ($\|\partial\phi_{t}/\partial t\|$ and NDV drastically decrease) while maintaining similar interpolation accuracy. Furthermore, the extrapolation similarity and flow consistency losses enable TimeFlow to predict plausible extrapolated deformation fields for future image estimation. These results demonstrate that TimeFlow inherently learns temporally continuous and spatially smooth deformation fields without explicit regularization, thus leading to the least trade-off between accuracy and smoothness. Visualization of the temporal deformations can be referred to Appendix.
6 Discussion and Conclusion
Novelty.
TimeFlow introduces novel interpolation and extrapolation constraints that ensure temporal consistency without temporal smoothness constraints. Together with the time-conditioning mechanism inspired by diffusion models, this enables effective extrapolation beyond the observed time frame. While we have used UNet as the base model, our contributions are generic and can, therefore, also be integrated into other registration backbones. Importantly, TimeFlow has significant application potential in areas such as in-silico drug effectiveness analysis. By extrapolating brain images to future timepoints at the beginning of a treatment, the predicted outcomes can be compared with actual follow-up scans. Discrepancies between predicted and observed results may then be attributed to treatment effects, offering a novel and efficient approach for evaluating therapeutic efficacy in clinical settings.
Limitation and failure cases.
TimeFlow encounters challenges in the following two scenarios: 1. Future image prediction/aging progression analysis with minimal biological aging differences: When the two base input images are temporally close and exhibit no noticeable aging differences, TimeFlow struggles to predict future images effectively based on them. In such cases, the two input images collapse into a single timepoint, rendering the model incapable of extrapolation based on a “single” input data. 2. Prediction for large time intervals: Although TimeFlow is designed to estimate future images for any given time parameter $t\in[0,\infty]$, its performance deteriorates for large $t$ values e.g. $t$ > 6. This limitation is expected due to the insufficient availability of long-range temporal data in the ADNI dataset, meaning TimeFlow rarely encountered such scenarios during training, and, of course, the fact that modeling such long developments, which are influenced by a plethora of parameters, is inherently very challenging.
Conclusion.
In this work, we introduced TimeFlow, a novel framework for longitudinal brain MR registration that addresses key challenges in the field, including the trade-off between accuracy and temporal consistency, limited sequential data, and the inability to predict future brain states. TimeFlow leverages a UNet-based architecture with temporal condition embeddings to ensure temporally continuous and spatially smooth deformation fields, eliminating the need for explicit smoothness constraints. Unlike existing methods that require full sequential data, TimeFlow learns directly from even just two images per patient, achieving both high registration accuracy and robust extrapolation to future timepoints. Our results show that TimeFlow outperforms baselines in both interpolation and extrapolation tasks and facilitates meaningful brain aging and neurodegenerative disease progression analysis. By bridging interpolation and future state prediction, TimeFlow facilitates clinically relevant analyses without extra anatomical annotations, representing a significant step toward data-driven models for tracking, predicting, and ultimately intervening in neurodegenerative diseases.
Acknowledgements
This work was supported by the German Federal Ministry for Economic Affairs and Climate Action (BMWi) through the project NeuroTEMP, and by the Munich Center of Machine Learning (MCML). Jiazhen Pan is funded by the European Research Council (ERC) grant Deep4MI (No. 884622).
The authors declare that they have no competing interests.
Supplementary Material
7 Pairwise Registration
Pairwise Baseline Endpoint
[…]
rigid 11.0$\pm$4.5 16.6$\pm$2.2 - -
Seq-VXM 7.0$\pm$4.1 21.2$\pm$2.9 3.4$\pm$0.8 0.8$\pm$2.7
NODER 8.0$\pm$3.9 20.0$\pm$2.6 3.2$\pm$1.3 18.8$\pm$89.6
VXM 7.3$\pm$4.5 20.9$\pm$2.9 3.0$\pm$0.6 15.2$\pm$73.2
VXMdiff 6.7$\pm$3.7 21.6$\pm$2.9 4.3$\pm$1.3 33.9$\pm$175.8
TimeFlow 7.3$\pm$4.5 20.8$\pm$2.9 3.2$\pm$0.2 2.0$\pm$4.1
TimeFlowdiff 6.6$\pm$4.2 21.7$\pm$3.0 4.5$\pm$0.5 0.3$\pm$0.5
Table 3: Quantitative registration results for Pairwise Registration between baseline and endpoint. Accuracy (foreground): MAE, PSNR; Smoothness: SDlogJ, NDV.
8 Temporal Continuous Deformation Fields
Figure 6: Visualization of temporal deformations $\phi_{t}$. $t$ is sampled from $t=0.5$ to $t=2.5$ with an interval $\Delta t=0.75$. The warped moving images are shown in Fig. 7.
Figure 7: Visualization of warped moving Images $I_{m}\circ\phi$ by the temporal deforamtons. The corresponding flow fields are shown in Fig. 6.
Figure 8: Visualization of temporal deformations $\phi_{t}$. $t$ is sampled from $t=0.5$ to $t=2.5$ with an interval $\Delta t=0.75$. The warped moving images are shown in Fig. 9.
Figure 9: Visualization of warped moving Images $I_{m}\circ\phi$ by the temporal deforamtons. The corresponding flow fields are shown in Fig. 8.
Figure 10: Visualization of temporal deformations $\phi_{t}$. $t$ is sampled from $t=0.5$ to $t=2.5$ with an interval $\Delta t=0.75$. The warped moving images are shown in Fig. 11.
Figure 11: Visualization of warped moving Images $I_{m}\circ\phi$ by the temporal deforamtons. The corresponding flow fields are shown in Fig. 10.