### v3

# IEEEexample:BSTcontrol

Abstract
Longitudinal brain analysis is essential for understanding healthy aging and identifying pathological deviations. Longitudinal registration of sequential brain MRI underpins such analyses. However, existing methods are limited by reliance on densely sampled time series, a trade-off between accuracy and temporal smoothness, and an inability to prospectively forecast future brain states. To overcome these challenges, we introduce TimeFlow, a learning-based framework for longitudinal brain MRI registration. TimeFlow uses a U-Net backbone with temporal conditioning to model neuroanatomy as a continuous function of age. Given only two scans from an individual, TimeFlow estimates accurate and temporally coherent deformation fields, enabling non-linear extrapolation to predict future brain states. This is achieved by our proposed inter-/extrapolation consistency constraints applied to both the deformation fields and deformed images. Remarkably, these constraints preserve temporal consistency and continuity without requiring explicit smoothness regularizers or densely sampled sequential data. Extensive experiments demonstrate that TimeFlow outperforms state-of-the-art methods in terms of both future timepoint forecasting and registration accuracy. Moreover, TimeFlow supports novel biological brain aging analyses by differentiating neurodegenerative trajectories from normal aging without requiring segmentation, thereby eliminating the need for labor-intensive annotations and mitigating segmentation inconsistency. TimeFlow offers an accurate, data-efficient, and annotation-free framework for longitudinal analysis of brain aging and chronic diseases, capable of forecasting brain changes beyond the observed study period.
{IEEEkeywords}
Brain MRI Image Registration, Longitudinal Analysis, Brain Aging Analysis
1 Introduction
Understanding how the human brain evolves throughout the lifespan, whether during healthy aging or in the presence of neuropathological disorders such as neurodegenerative diseases, remains a central challenge in neurology and neuroradiology [1, 2]. Distinguishing these trajectories not only clarifies pathological mechanisms but also refines prognostic models, ultimately supporting timely and personalized interventions [3]. Magnetic Resonance Imaging (MRI) has become the standard tool for capturing longitudinal brain changes, owing to its high spatial resolution and superior soft-tissue contrast. Repeated imaging sessions allow clinicians and researchers to investigate evolving patterns of atrophy, lesion progression, and other structural alterations over time [4, 5].
Figure 1: Top: Given two longitudinal scans (visits at Baseline and Year 4 visits), TimeFlow can interpolate and extrapolate novel brain MRI scans at unseen timepoints (e.g., Year 2 and Year 7). Bottom: Visual comparison of extrapolated predictions at Year 7 demonstrates TimeFlow’s superior performance, compared to (i) NODER, which uses NeuralODE-based geodesic regression, and (ii) VoxelMorph, which linearly scales displacement field. Prediction errors are marked by red arrows.
At the core of most quantitative longitudinal brain analyses lies longitudinal registration, which aligns a subject’s brain scans across multiple timepoints [6, 7]. Despite steady advances in longitudinal registration, two critical challenges persist: 1. Ensuring temporal continuity with limited sequential data. Existing methods often rely on explicit temporal smoothness constraints [8, 9, 10] which usually compromise accuracy. Besides, these methods often require densely sampled longitudinal data (often more than five scans) [11, 12], which is impractical in typical clinical scenarios where usually fewer than three follow-up scans are available. 2. Predicting future brain states and analyzing aging progression. While retrospective assessments of past changes are informative, prospective forecasting of brain anatomy would markedly enhance clinical utility [13]. Such forecasting capabilities enable clinicians to anticipate accelerated disease progression and adjust interventions proactively [14]. However, most recent methods model only the observed scan sequence [15, 16, 17], allowing interpolation between existing timepoints. It limits their capability in extrapolating beyond existing scans, as illustrated in Figure˜1, thereby restricting their ability to assess deviations from expected progression, anticipate disease worsening, and adjust treatment accordingly.
To overcome these limitations, we introduce TimeFlow, a novel learning-based longitudinal brain MRI registration framework. TimeFlow combines a U-Net backbone with temporal conditioning to model neuroanatomy as a continuous function of age. As shown in Figure˜2, TimeFlow requires only two scans per subject to estimate the bidirectional deformation fields for arbitrary timepoints, facilitating both interpolation (timepoints within observed intervals) and extrapolation (timepoints beyond observed intervals). Our proposed bidirectional inter-/extra-polation consistency constraints allow unlimited sampling of timepoints, rather than the limited discrete visits during training, ensuring temporally coherent deformation fields. This synergy between the temporal conditioning and the proposed constraints enables TimeFlow to predict future brain states and supports aging analyses without additional annotations. Unlike generative models that create synthetic MRI scans, TimeFlow uses image warping to ensure anatomically plausible outputs, maintain consistency, and provide interpretability through deformation field inspection.
Our contributions are threefold:
  1. 1.
    Temporally-conditioned registration. A unified network conditioned on a continuous time parameter produces temporally coherent deformation fields from only two input scans, accurately predicting unobserved anatomical states.
  2. 2.
    Bidirectional inter-/extra-polation consistency loss. Symmetric supervision on bidirectional deformations removes the need for explicit temporal smoothness regularizers, significantly improving temporal coherence, continuity, and generalization through unlimitedly sampled timepoints.
  3. 3.
    Registration-only aging analysis. A registration-only pipeline quantifies relative aging trajectories by referencing healthy controls, differentiating Mild Cognitive Impairment (MCI) and Dementia without requiring tissue labels, thus removing the need for non-trivial annotations.
Figure 2: TimeFlow takes source and target images as input, conditioned on the continuous temporal parameter $t$. Adjusting $t$ enables the generation of interpolated brain scans within the observed interval ($0<t<1$), or extrapolated scans for future ($t>1$).
2 Related Work
2.1 Registration across subjects
Cross-sectional registration methods laid the groundwork for medical image registration, pioneering this field with techniques such as parametric spline-based models [18, 19], optical flow [20, 21], and diffeomorphic transformations [22, 23, 24]. However, they demand prohibitive computing time in the order of several hours to register one single subject, posing a significant barrier to clinical deployment. To address this, learning-based frameworks [25, 26, 27, 28, 29, 30] have emerged, dramatically reducing inference time and often improving registration accuracy. Nevertheless, all such methods rely on carefully tuned smoothness regularizers to ensure anatomically plausible deformation fields. An improper weighting factor of smoothness regularizers can either reduce fidelity or introduce unrealistic warps. Additional tuning or adaptive strategies are often needed to strike an optimal balance [31, 32, 33].
2.2 Registration across time
Longitudinal registration methods [34, 35, 36, 37, 38, 12, 10, 39] often leverage the backbone architecture of the cross-sectional registration, whereas one additional constraint is introduced: the temporal continuity/consistency across multiple timepoints of a subject. Two major strategies are often used to tackle this problem. First, an explicit smoothness constraint along the temporal dimension is incorporated to maintain coherent deformation fields over time [9, 11]. However, as with spatial smoothness, temporal regularization faces a similar trade-off: enforcing temporal continuity can inadvertently lower spatial accuracy. Second, many longitudinal methods require multiple, densely sampled scans (>10) to capture temporal dynamics [16, 15]. In real-world clinical scenarios, patients often have only a few follow-up scans ($\leq 3$), restricting model flexibility and practical implementation. Importantly, these methods commonly can only carry out retrospective analyses and are not capable of prospective studies, such as future image extrapolation and aging/disease progression risk prediction.
2.3 Longitudinal brain analysis
Existing longitudinal studies of neurodegenerative diseases like Alzheimer’s disease usually require additional annotations, such as cortical surfaces and structural segmentations to quantify the brain changes. However, these labels are usually created by cross-sectional pipelines that process each visit independently, which are both time-consuming and prone to temporal inconsistency, i.e., the same anatomy receives different labels at successive visits. Consequently, quantities derived from those labels (e.g., hippocampal volume, cortical-thickness trajectories) require post-processing to suppress the systematic noise [40, 41]. These limitations motivate annotation-free, temporally coherent models that treat the sequential series as a single spatiotemporal object instead of a collection of disconnected timepoints.
Figure 3: Left: Architecture of TimeFlow. The network $f_{\theta}$ incorporates temporal conditioning via adaptive instance normalization. It takes two input images ($I_{0}$$I_{L}$) and a continuous time variable $t$ to predict the temporal deformation field $\phi_{t}^{0\rightarrow L}$. Examples of $t=0.5$ (interpolation at middle), $t=1$ (endpoint registration), and $t=2$ (extrapolation) are shown. Right: Interpolation (upper yellow) and Extrapolation (bottom green) Consistency Constraints. The symmetric intermediate image similarity enforces the deformed source and target to be aligned at time $t_{k}$, while the interpolation flow consistency loss encourages the deformation composition consistency. The extrapolation similarity and flow constraints give information on what a realistic and reasonable forecast should be. Training with both consistency losses enables TimeFlow to generate plausible deformation in $t\in[-\infty,+\infty]$. The input images are highlighted by blue squares (Input), while the supervision is highlighted by dashed squares (Sup.).
2.4 Future prediction via generative models
Recent advances in diffusion-based generative models have inspired several attempts to synthesize longitudinal brain MRIs conditioned on variables such as age and diagnosis [42, 43]. Despite producing visually plausible synthetic images, these methods largely remain black-box, as their diffusion trajectory does not map clearly onto biophysical deformations. This limits the interpretability, particularly for clinicians interested in tracking local changes like region-specific atrophy. Besides, diffusion models typically require large-scale datasets for stable training, which is impractical in clinical settings where follow-up is sparse. Nonetheless, generative models retain promise as complementary tools in neuropathological disorders involving intensity anomalies, such as gliomas or lesions, where longitudinal registration alone is insufficient to represent newly emerging tissue structures.
3 Methods
3.1 Time-Continuous Registration
Given a longitudinal MRI sequence $\{I_{k}\mid k=0,\ldots,L\}$, our goal is to derive a temporally continuous deformation trajectory connecting all intermediate timepoints. To achieve this, we propose a time-continuous registration framework, TimeFlow, defined as $\phi_{t}=f_{\theta}(I_{0},I_{L},t)$. As illustrated in Figure˜2, TimeFlow accepts the baseline image $I_{0}$ (source) and the last image $I_{L}$ (target) and predicts the deformation fields $\phi_{t}$ parametrized by continuous time $t$. For interpolation within the observed interval ($0<t<1$), TimeFlow predicts intermediate states between source and target scans. For extrapolation ($t>1$), TimeFlow predicts future brain states beyond the last observed visit. To enforce temporal coherence and continuity in the deformation trajectories, we introduce the novel inter-/extra-polation consistency constraints. The core architecture (Figure˜3, Left) is a time-embedded U-Net [44, 45]. Specifically, a sinusoidal positional encoding and multi-layer perceptrons (MLPs) project the scalar time variable $t$ into a latent embedding space. Next, the time latent vector is fed to two separate MLPs to steer the mean and variance of the inter-layer adaptive instance normalization. The network is thus instilled with the capability of generating meaningful time embedding and predicting time-continuous flow fields by the novel time conditioning mechanism.
3.2 Interpolation Consistency
Symmetric Intermediate Similarity
We first define the time-continuous registration function as $\phi_{t}^{0\rightarrow L}=f_{\theta}(I_{0},I_{L},t)$, where $f_{\theta}(I_{0},I_{L},0)=\textrm{Id}$ is the identity mapping at $t=0$. When $t\in(0,1]$, $\phi_{t}^{0\rightarrow L}$ is the forward deformation mapping from $I_{0}$ towards $I_{L}$, and $\phi_{1}^{0\rightarrow L}$ moves $I_{0}$ to reach $I_{L}$. Similarly, for negative values, $t<0$, $\phi_{t}$ maps backward from $I_{L}$ to $I_{0}$. For any intermediate image $I_{k}$ observed at time $t_{k}$, we can formulate the symmetric interpolation image similarity measure as:
$\mathcal{L}_{\textrm{sim}}(I_{0}\circ\phi_{t_{k}}^{0\rightarrow L},I_{k})+\mathcal{L}_{\textrm{sim}}(I_{L}\circ\phi_{t_{k}-1}^{0\rightarrow L},I_{k}),$ (1)
which deforms both source and target images to remove the bias introduced by single-side warping [46, 47, 48]. As shown in Figure˜3 (Right, top yellow), $I_{0}\circ\phi_{t_{k}}^{0\rightarrow L}$ is warping $I_{0}$ to the time $t=t_{k}$ and $I_{L}\circ\phi_{t_{k}-1}^{0\rightarrow L}$ is warping $I_{L}$ to the time $t=t_{k}$. This symmetry enforces bidirectional consistency, facilitating plausible deformation trajectories passing through observed intermediate scans.
Generalized Symmetric Intermediate Similarity
However, in real-world clinical scenarios, many subjects have fewer than three MRI scans available. In such a case, we generalize the above constraint by synthetically sampling intermediate timepoints $\hat{t}$ from a uniform distribution $\mathcal{U}(0,1)$, defining the generalized interpolation similarity as:
$\mathcal{L}_{\textrm{sim}}(I_{0}\circ\phi_{\hat{t}}^{0\rightarrow L},I_{L}\circ\phi_{\hat{t}-1}^{0\rightarrow L}),$ (2)
Here, both warped images $I_{0}\circ\phi_{\hat{t}}^{0\rightarrow L}$ and $I_{L}\circ\phi_{\hat{t}-1}^{0\rightarrow L}$ serve as a synthetic version of intermediate timpoint $I_{\hat{t}}$, eliminating reliance on actual intermediate observations. By freely sampling $\hat{t}$, TimeFlow effectively learns densely and continuously across the entire temporal axis.
Interpolation Flow Consistency
To further encourage plausible time-continuous deformations, we propose the interpolation flow consistency as a regularizer:
$\begin{split}\mathcal{L}_{\textrm{inter-consis-flow}}=&\|\phi_{1}^{0\rightarrow L}\circ\phi_{t_{k}-1}^{0\rightarrow L}-\phi_{t_{k}}^{0\rightarrow L}\|_{2}^{2}\\ +&\|\phi_{-1}^{0\rightarrow L}\circ\phi_{t_{k}}^{0\rightarrow L}-\phi_{t_{k}-1}^{0\rightarrow L}\|_{2}^{2},\end{split}$ (3)
which combines the two constraints shown in Figure˜3 (Right, top yellow). We formulated the loss with the notation $t_{k}$, but as described in the previous paragraph, $t_{k}$ can also be sampled from the uniform distribution if no intermediate observation exists. It is derived from the common transformation composition consistency that two consecutive transforms are equivalent to a direct transform. Moving from $t=0$ to $t=1$ and then back to $t=t_{k}$ is equivalent to moving directly from $t=0$ to $t=t_{k}$ (arrow 1 composed with arrow 3 should equal to arrow 2 in Figure˜3, Right). The backward case is derived similarly. It is worth noting that this differs from the commonly used smoothness regularizations, which explicitly enforce a smooth gradient over the flow field. Therefore, there is only a minimal trade-off between accuracy and smoothness since both the symmetric image similarity and the flow consistency encourage plausible and reasonable longitudinal transformations.
3.3 Extrapolation Consistency
Through the intermediate similarity and interpolation flow consistency, we obtain meaningful predictions for bidirectional registrations in $t\in[-1,1]$. Now we introduce extrapolation constraints for forward ($t>1$) and backward (t<1t<-1) directions. To this end, we consider the deformation $\phi_{1}^{0\rightarrow k}$ from $I_{0}$ to $I_{k}$. Given the time of $I_{k}$ being $t_{k}$, it holds that $\phi_{t_{k}}^{0\rightarrow L}=\phi_{1}^{0\rightarrow k}$. As illustrated in Figure˜3 (Right), they both register $I_{0}$ to $I_{k}$. Consequently, we can then formulate $\phi_{1}^{0\rightarrow L}=\phi_{1/t_{k}}^{0\rightarrow k}$ (arrow 1 equals to arrow 5 in Figure˜3, Right), where $\phi_{1/t_{k}}^{0\rightarrow k}$ denotes an extrapolation because $t_{k}\in(0,1)$ and $1/t_{k}>1$. Based on this notation, we define the extrapolation image similarity as:
$\mathcal{L}_{\textrm{sim}}(I_{0}\circ\phi_{1/t_{k}}^{0\rightarrow k},I_{L}).$ (4)
Again, if no intermediate observations are available, we can sample $t_{k}$ from $\mathcal{U}(0,1)$ and define $I_{k}=I_{0}\circ\phi_{t_{k}}^{0\rightarrow L}$ to create synthetic intermediate data. In addition to the extrapolation similarity measure, we define the forward extrapolation flow consistency as:
$\mathcal{L}_{\textrm{ext-consis-flow}}=\|\phi_{1}^{0\rightarrow L}-\phi_{1/t_{k}}^{0\rightarrow k}\|_{2}^{2}.$ (5)
As shown in Figure˜3 (Right), both $\phi_{1}^{0\rightarrow L}$ (arrow 1) and $\phi_{1/t_{k}}^{0\rightarrow k}$ (arrow 5) map from $I_{0}$ to $I_{L}$, thus should be consistent with each other. The backward versions of the extrapolation similarity and extrapolation flow consistency are similarly defined as:
$\begin{split}&\mathcal{L}_{\textrm{sim}}(I_{L}\circ\phi_{1/(t_{k}-1)}^{k\rightarrow L},I_{0})\quad\text{and}\\ &\mathcal{L}_{\textrm{ext-consis-flow}}=\|\phi_{-1}^{0\rightarrow L}-\phi_{1/(t_{k}-1)}^{k\rightarrow L}\|_{2}^{2}.\end{split}$ (6)
The similarity measures and flow consistencies both contribute to achieving a realistic and coherent extrapolation forecasting. The proposed losses effectively ensure accurate extrapolation while implicitly capturing temporal consistency and continuity, without the need for additional temporal smoothness constraints.
3.4 Unlimited Sampled Triplets
As illustrated in Figure˜3 (Right), the proposed inter-/extra-polation constraints require observed triplets of timepoints, limiting learning when few discrete scans are available. To address this, we use “synthetic” triplets dynamically sampled from continuous time intervals. Specifically, by sampling intermediate timepoints $\hat{t}$ from $\mathcal{U}(0,1)$ during training, we can generate infinite triplets from a pair of source and target images, freeing TimeFlow from the discrete, fixed timepoints. This continuous, densely sampled temporal supervision significantly enhances the network’s capacity to generalize and accurately model brain anatomy changes at any arbitrary timepoint, including unseen future scenarios, while using only two scans as inputs.
4 Data and Experiments
Data and Pre-processing
We train and evaluate TimeFlow on the public Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset [4], which comprises longitudinal T1-weighted Magnetic Resonance (MR) scans from both healthy subjects and patients diagnosed with Alzheimer’s Disease (AD) or Mild Cognitive Impairment (MCI), i.e., neurodegenerative disorders that exhibit specific patterns of atrophy development in the brain. We employ the longitudinal processing from FreeSurfer v7.2 [49]. First, all scans are processed individually, including skull stripping, normalization. Subsequently, the image sequence is rigidly aligned via a groupwise rigid registration [50]. FreeSurfer aggregates information across time points and refines the pre-processed outputs. To ensure meaningful longitudinal analysis, we select 134 subjects exhibiting substantial structural changes (median brain deformation >0.3>0.3mm). Then, the dataset is divided into training, validation, and testing sets, ensuring a balance in diagnosis, age, sex, and the number of visits. It results in 74, 30, 30 subjects for each split, respectively, with the number of time points per subject ranging from 2 to 12. All scans have isotropic 1mm spacing and are cropped to a uniform size of $160\times 160\times 192$ voxels. The intensity values are scaled to the range [0,1], from percentile [0,99.9] without clipping. Additionally, 11 subjects from the OASIS dataset [51] serve for zero-shot evaluation.
Figure 4: Visualization of extrapolation errors compared to ground truth future images. Coronal and axial views of two subjects are presented. Areas with high error are marked by red arrows.
Implementation Details
Our model is implemented using Pytorch 2.1.2. and training/test is carried out on one NVIDIA A100 40G with a batch size of $2$ (sample $\hat{t}$ twice per subject). We set the dimension of time embedding to $16$, the feature channels in the U-Net encoder and decoder to $[32,32,48,48,96]$. In addition to TimeFlow, which outputs the time-dependent deformation field directly, we have the diffeomorphic variant TimeFlowdiff, which utilizes the scaling and squaring layer with $7$ integration steps to integrate the temporal stationary velocity field. The weights of both interpolation and extrapolation similarity are set to $1.0$. The weights of interpolation and extrapolation flow consistency are set to $2,0.03$ for TimeFlow and $1.25,0.025$ for TimeFlowdiff.
Figure 5: Future image predictions at different extrapolation time $t$, where no ground truth is available. $t=1$ represents one-step pairwise registration between $I_{m}$ and $I_{f}$. Thus, $t=2,4,6$ represents two-, four-, and six-step extrapolation over $I_{f}$. Major errors are highlighted with red arrows. VXM/diff produces unsmooth deformations, leading to structural inconsistencies and loss of anatomical details (zoom in for more details).
[…]
Extrapolation Interpolation
[…]
(Upper bound)
[…]
Figure 6: Visualization of interpolation (Year 1.5) and extrapolation (Year 8) results from coronal and axial views of a subject’s brain. TimeFlow/diff takes the Baseline and Year 4 images as input, interpolates to Year 1.5, and extrapolates to Year 8. Ground truth images (row 1) and warped images (row 2) are compared. Initial misalignment (row 3) is effectively corrected (row 4), demonstrating TimeFlow’s accuracy in temporal modeling. Error values range from [1,1][-1,1] due to the image intensity normalization to $[0,1]$.
Baselines
To evaluate the registration performance, we compare TimeFlow with the following methods:
  • NODER [17, 16] takes the complete image sequence as the model input and is thus aware of every time point. It carries out subject-specific optimization and models the time-varying velocity field with a neural network and leverages NeuralODE [52] to compute integration and adjoint gradients. It fits the subject-specific temporal deformation trajectory by maximizing the similarity with each frame.
  • VoxelMorph (VXM) [25] and the diffeomorphic variant (VXMdiff[53]. Since both are pairwise registration methods, the interpolation and extrapolation prediction is carried out based on the linear scaling of the displacement or velocity field from the pairwise registration. The number of integration steps of VXMdiff is also set to 7.
  • Sequence-based VoxelMorph (Seq-VXM) [15]. Seq-VXM processes the whole sequence and registers the baseline image to each intermediate time point. Given that Seq-VXM performs multiple pairwise registrations by taking the full sequence as model input, we treat Seq-VXM as the performance upper bound for the retrospective study (refer to Tab. 1).
All methods use Local Normalized Cross Correlation (LNCC) with radius $r=9$ (NODER) or $r=4$ (others) as image similarity loss. The loss function weighting factors in NODER follow their original suggestion. Seq-VXM, VXM and VXMdiff compute the SOTA smoothness regularizer GradICON [54, 55] with weight $1.5,0.5,1.0$.
Figure 7: Left: Prospective, annotation-free analysis of brain aging progression rates for Control (CN), MCI and Dementia. Rates near 1 reflect normal, chronological brain aging, whereas values above 1 indicate accelerated neurodegenerative progression. Right: Retrospective group-wise analysis of longitudinal brain progression. Trajectories represent median linear fits between TimeFlow’s inferred temporal brain state ($t$) and chronological years since baseline. Shaded ribbons indicate the 40th-60th percentile range. A steeper slope corresponds to faster neuroanatomical changes. Colormap: CN-green, MCI-orange, Dementia-red.
Figure 8: Quantitative comparison of extrapolation accuracy among TimeFlow and baseline methods on ADNI dataset. Barplots represent mean values with standard errors. Significance levels by Wilcoxon $t$-test (Holm–Bonferroni correction) are indicated. ***:p<0.001, **:p<0.01, *:p<0.05
Evaluation metrics
[…]
5 Results
5.1 Prospective studies: Predicting the future
We first evaluate TimeFlow’s performance in prospective scenarios, specifically, its ability to extrapolate deformation fields to predict future brain scans. TimeFlow supports two forms of prospective analysis: (1) Given two available scans and a future time variable $t>1$, the model directly predicts the corresponding deformation fields and resulting future MRI images; (2) Given an observed future scan, TimeFlow can infer the corresponding $\phi_{t}$ and use the estimated time variable $t$ as a surrogate metric to quantify the biological aging progression.
Given t, predict the future image
A significant advantage of TimeFlow, as described in Section˜3, is its ability to extrapolate deformation fields to future timepoints beyond the observed interval. The quantitative performance of TimeFlow’s extrapolation forecasting is reported in Table˜1. The metrics are computed with respect to the ground truth future scans. Results clearly demonstrate that TimeFlow consistently achieves superior extrapolation accuracy compared to baseline methods, yielding the lowest MAE and highest PSNR (Figure˜8), even when evaluated in a zero-shot manner on the unseen OASIS dataset. Note that Seq-VXM cannot carry out future predictions; thus, it is excluded from this comparison. Qualitative extrapolation results are visualized in Figure˜6 (Year 8) and Figure˜4, illustrating TimeFlow’s accuracy by comparing to the ground truth images. Moreover, Figure˜5 demonstrates extrapolation at a large $t$ ($t=2,4,6$) where ground truth is not available. In these cases, NODER produces unrealistic and anatomically implausible deformations, resulting in significant errors in future image prediction. Similarly, VXM/diff, which applies linear scaling to displacement/velocity fields, generates unsmooth deformations (indicated by high NDV) and dissolves the brain, causing the anatomical details to fade. In contrast, TimeFlow/diff consistently generates smooth and coherent extrapolated deformation fields, quantitatively supported by the lowest MAE and minimal NDV values in Table˜1.
Given future scan, assess the brain aging
As emphasized in Section˜1, the extrapolation capability of TimeFlow enables annotation-free disease progression analysis. We compare the brain aging rates across diagnostic groups, i.e., Controls (CN), MCI, and Dementia, by predicting the optimal time parameter ($t_{\textrm{ext}}$) corresponding to the observed future MRI scans. Specifically, MRI triplets are sampled from validation and test subsets, where the intervals between the first and second visits match closely the intervals between the second and third visits. The diagnosis at the third visit labels each triplet. Using the first two scans as input, TimeFlow extrapolates forward to the third, future scan. By comparing predictions against actual observed future, we compute the extrapolated time parameter $t_{\textrm{ext}}$, and divide it by the normalized chronological intervals (in years) to yield a brain aging rate. An aging rate above 1 indicates accelerated pathological changes, as expected in patients with neurodegenerative conditions. Indeed, the boxplot of Figure˜7 (Left) clearly shows that Dementia patients exhibit the highest progression rate, followed by MCI subjects, while the control group maintains progression rates around 1. The analysis validates that our annotation-free surrogate brain aging rate metric is able to separate disease groups.
[…]
$\mathcal{L}_{\textrm{inter}}$ $\mathcal{L}_{\textrm{ext}}$ Extrapolation Interpolation
[…]
5.2 Retrospective studies: Matching the deformation trajectory
We further evaluate TimeFlow’s interpolation capability, assessing its effectiveness in modeling temporally continuous and anatomically plausible deformation trajectories connecting observed MRI visits. In this setting, TimeFlow uses only the baseline visit $I_{0}$ and the final visit $I_{L}$ as inputs, predicting deformation fields $\phi_{t_{k}}$ for intermediate images $I_{k}$, where $0<k<L$. Quantitative comparisons with baseline methods for interpolation are presented in Table˜1. As discussed in baselines, Seq-VXM, performing explicit pairwise registrations for every available timepoint, represents a performance upper bound. Despite using only the baseline and last scan as inputs, TimeFlow outperforms NODER on the ADNI dataset, which is also aware of the entire sequence and parametrized by Neuralordinary differential equation (ODE). TimeFlow achieves comparable performance to NODER in zero-shot evaluation on the OASIS dataset, where NODER performs subject-specific optimization on the full image sequence. Additionally, TimeFlowdiff achieves the least non-diffeomorphic volume without extra spatial smoothness regularization. Qualitative visualizations in Figure˜6 (Year 1.5) illustrate the accuracy of TimeFlow’s interpolations, substantially reducing alignment errors compared to initial misalignments. Similar to our above prospective analysis, we also perform retrospective, group-wise brain progression analysis using the inferred temporal parameter $t$ of intermediate visits as a surrogate brain progression state marker. We extract trajectory segments with the same diagnoses from each subject, and then compute a linear relationship between the estimated brain state ($t$) and chronological age (years since baseline of the segment). The median linear lines of the groups (CN, MCI, Dementia) in Figure˜7 (Right) clearly illustrate distinct progression patterns, with Dementia exhibiting the fastest structural changes, MCI intermediate, and CN the slowest.
5.3 Ablation Study
As described in Section˜3 and supported by both quantitative and qualitative results, TimeFlow achieves temporally continuous and spatially smooth deformation fields without requiring explicit smoothness regularization. To assess the effectiveness of our proposed inter-/extra-polation consistency constraints, we perform detailed ablation studies. To evaluate the temporal smoothness and continuity of TimeFlow, we compute the derivative $\partial\phi_{t}/\partial t$ using the finite difference method by sampling $t$ at intervals of $\Delta t=0.25$. Similar to the commonly used diffusion regularizer in learning-based registration methods, which computes the Dirichlet energy over spatial gradients, we quantify temporal smoothness as $\|\partial\phi_{t}/\partial t\|$. Specifically, we compute the norm of the temporal displacement derivatives within the brain regions and then take the mean over all brain voxels as the final metric. As shown in table˜2, incorporating interpolation consistency loss significantly improves both temporal and spatial smoothness ($\|\partial\phi_{t}/\partial t\|$ and NDV drastically reduce) without sacrificing interpolation accuracy. Furthermore, the extrapolation similarity and flow consistency losses enable TimeFlow to generate plausible extrapolated deformation fields for future image estimation. These ablation results demonstrate that TimeFlow inherently learns temporally continuous and spatially smooth deformation fields, achieving an optimal balance between registration accuracy and smoothness without explicit smoothness penalties.
6 Discussion
Contributions
TimeFlow introduces novel interpolation and extrapolation constraints that ensure temporal consistency without explicit temporal smoothness regularization. Combined with the time-conditioning mechanism, TimeFlow enables effective extrapolation beyond the observed time frame. While we have used U-Net as the base model, our contributions are generic and can be integrated into other registration backbones. Importantly, TimeFlow has promising application potential in areas such as in silico drug effectiveness analysis. By extrapolating brain images to future timepoints at the beginning of a treatment, the predicted outcomes can be compared with actual follow-up scans. Discrepancies between predicted and observed results could then be attributed to treatment effects, offering a novel and efficient approach for evaluating therapeutic efficacy in clinical settings.
Bidirectional extrapolation
While we mainly focus on the forward extrapolation (predicting future brain states) in this study, TimeFlow is inherently bidirectional and allows backward extrapolation as well. Backward extrapolation, i.e., traveling back in time and visualizing the brain at earlier, unobserved ages, opens several avenues: (i) reconstructing a pre-symptomatic baseline for patients before neurodegeneration; (ii) generating counterfactual controls that share a subject’s anatomy but simulate a healthy aging trajectory.
Limitation and failure cases
TimeFlow encounters challenges in the following two scenarios: 1. Future image prediction/aging progression analysis with minimal biological aging differences: When the two input scans are temporally close and exhibit no noticeable structural differences, they collapse into a single timepoint, which TimeFlow struggles to predict future images effectively. 2. Prediction for far future: Although TimeFlow is designed to estimate future images for any given time parameter $t\in[0,\infty]$, its performance deteriorates for large $t$ values, e.g., $t$ > 6. This limitation is expected due to limited data covering such long-range time intervals in the ADNI dataset, meaning TimeFlow rarely encountered such scenarios during training. The modeling of such long developments, which are influenced by a plethora of parameters, is inherently challenging.
7 Conclusion
We introduced TimeFlow, a novel framework for longitudinal brain MR registration, specifically designed to address persistent challenges such as limited sequential data, the accuracy-temporal consistency trade-off, and the inability to predict future brain states. TimeFlow leverages a U-Net-based architecture with temporal condition embeddings to ensure temporally continuous and spatially smooth deformation fields, eliminating the need for explicit smoothness penalties. Unlike existing methods that require full sequential data, TimeFlow learns effectively from only two scans per patient, achieving both high registration accuracy and robust extrapolation to future timepoints. Extensive evaluations demonstrate TimeFlow’s superior performance over existing baselines in both interpolation and extrapolation tasks, as well as its ability to facilitate clinically meaningful analyses of brain aging and neurodegenerative diseases. By seamlessly bridging retrospective interpolation and prospective extrapolation, TimeFlow enables annotation-free analyses, representing a significant step toward data-driven models for tracking and predicting neuroanatomical changes in neurodegenerative diseases.